# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

### v1

@@ 6.2 Leave-one-fold-out Feature Extraction
To address this challenge, we propose a leave-one-fold-out strategy that enables the extraction of comparable embeddings for training and test data. Within the TabPFN v2 framework, we define the support set ${\mathcal{S}}$ as the examples with true labels and the query set ${\mathcal{Q}}$ as the examples with dummy labels. Under the standard configuration, ${\mathcal{S}}$ corresponds to the labeled training set, while ${\mathcal{Q}}$ corresponds to the unlabeled test instances.
A key trade-off arises: ${\mathcal{S}}$ needs to include as much labeled data as possible to propagate knowledge from ${\mathcal{S}}$ to ${\mathcal{Q}}$ effectively in the in-context learning process. However, to ensure that embeddings for training and test instances are comparable, training instances must also be included in ${\mathcal{Q}}$ and paired with dummy label embeddings.
To reconcile these requirements, we split the training set into multiple folds (e.g., 10 folds). One fold is used as ${\mathcal{Q}}$ for embedding extraction, while the remaining folds constitute ${\mathcal{S}}$ with true training labels. This ensures that ${\mathcal{S}}$ retains sufficient label information while allowing embeddings to be extracted for the training data in ${\mathcal{Q}}$. An extreme version of this strategy operates in a leave-one-out fashion, with only one training instance placed in ${\mathcal{Q}}$ at a time.
Results in Figure 5 (c)-(f) demonstrate that embeddings extracted using this strategy (with 10 folds) more effectively reveal dataset properties. We observe that TabPFN v2 simplifies tabular data distributions and transforms datasets into nearly separable embedding spaces, particularly in the embeddings after intermediate transformer layers.
Table 2: Average rank (lower is better) of TabPFN v2 and a linear classifier trained on the extracted embeddings across 29 classification datasets. “Combined” refers to an approach where embeddings from up to three layers (from the 12 available layers) are selected based on the validation set performance and concatenated.
@@
Rank 3.12 3.43 5.03 4.72 2.45 2.24
6.3 Validation of Embedding Quality
To validate the quality of the extracted embeddings, we train a logistic regression on top of the extracted embeddings. Specifically, the classifier is trained on embeddings derived from the training set and evaluated on test set embeddings. The average rank across 29 classification datasets from the tiny benchmark2 in Ye et al. (2024b) is reported in Table 2.
Remarkably, training a linear classifier on these extracted embeddings yields performance comparable to TabPFN v2’s in-context learner. Embeddings from intermediate layers (or their selected concatenations) sometimes achieve even better results. These findings highlight TabPFN v2’s potential as a robust feature encoder, offering valuable insights into its architecture and paving the way for broader applications in tabular data analysis and representation learning.
7 Extensions of TabPFN v2
We aim to expand TabPFN v2’s applicability beyond the boundaries outlined in Hollmann et al. (2025), which restricts its use to datasets with no more than 10,000 samples, 500 dimensions, and 10 classes. To address these limitations without requiring extensive retraining, we propose post-hoc divide-and-conquer strategies inspired by Chain-of-Thought (CoT) prompting in large-language models (Wei et al., 2022). We reformulate challenging tasks into multiple simpler subtasks that TabPFN v2 can effectively handle.
Figure 6: Mean accuracy on 18 high-dimensional datasets. “*” denote our extension of TabPFN v2, and “-PCA” is another variant using PCA to reduce dimensions.
@@ 7.1 High Dimension Datasets
High-dimensional datasets (Jiang et al., 2024) present a unique challenge due to the quadratic complexity of TabPFN v2 with respect to the number of dimensions. To mitigate this, we propose subsampling the feature space into smaller subsets, processing each subset independently, and combining the predictions in an ensemble (bagging) fashion, similar to random forests (Breiman, 2001).
In detail, we iteratively sample $m$ subsets (we set $m=4$ in experiments), each containing $d^{\prime}<d$ randomly selected attributes. For each subset, we leverage the TabPFN v2’s ability to handle lower-dimensional data to obtain predictions. The final result aggregates these outputs using averaging (for regression) or majority voting (for classification).
Figure 6 summarizes the results 18 high-dimensional classification datasets. TabPFN v2 with the proposed divide-and-conquer and then ensemble strategy significantly increases the mean accuracy (the second-highest) against the vanilla TabPFN v2, effectively extending TabPFN v2’s scalability to datasets with $d\geq 2000$. Another variant that utilizes PCA to reduce the dimensionality together with bagging also resolves the dimensionality issue to some extent.
Figure 7: Mean accuracy on 12 datasets with more than 10 classes. “PFNv2” denotes our extension, while “PFNv2-ECOC” denotes the multi-class ECOC strategy implemented by Hollmann et al. (2025) in their February 2025 code update.
@@ 7.2 Multi-Class Problems with More Than 10 Classes
For a task with $C>10$ classes, we encode each label $y\in[C]$ as a $t$-digit decimal representation, where $t=\lceil\log_{10}C\rceil$. For each digit position $j\in\{1,\ldots,t\}$, we train a separate TabPFN v2 model $f_{j}$ to predict the $j$-th digit. For example, in a 15-class problem, we decompose it into two subtasks: the first predicts the tens digit (with classes ${0,1}$), while the second predicts the ones digit (with classes ${0,\dots,9}$). During inference, the predicted digits are reconstructed to obtain the final class label. To improve robustness, we permute the class order $\sqrt{C}$ times, leveraging the efficiency of TabPFN v2 to predict across all permutations. The final prediction is obtained by ensembling these results. As shown in Figure 7, this approach achieves the second-best mean accuracy on 12 datasets with more than 10 classes while preserving computational efficiency.
Figure 8: Average rank on 18 large-scale datasets. “*” indicates our extension. “-B” refers to the variant that randomly subsamples 10,000 training examples four times and aggregates their predictions. “-K” denotes the variant that selects a representative subset of 10,000 training examples based on proximity to prototypes obtained via KMeans. All variants improve TabPFN v2.
@@ 7.3 Large-Scale Datasets
For large-scale datasets, we randomly sample 10,000 training examples from the full training set as the support set and treat the remaining training examples and test instances as the query set. We extract their embeddings to form a new tabular dataset, on which a logistic regression classifier is trained to make predictions on the test set embeddings. This process is repeated four times, and the final predictions are aggregated, denoted as TabPFN v2-SQ.
We also investigate integrating TabPFN v2 with decision trees to handle large-scale tasks. Specifically, we first train a shallow decision tree, setting its maximum depth to 5 layers for datasets with fewer than 100,000 samples, and to 10 layers for larger datasets. The decision tree partitions the training set into smaller, more manageable subsets. Then, we apply TabPFN v2 to each leaf node, treating the examples in each node as a separate training set for TabPFN v2. During inference, a test instance is first passed through the shallow decision tree to a leaf node and then predicted by the corresponding TabPFN v2 model. We denote this extension as TabPFN v2-DT. We note that a similar strategy was mentioned in Hollmann et al. (2025) but for a drastically difference purpose. Specifically, Hollmann et al. (2025) use decision trees to handle within-dataset heterogeneous for some highly challenging tabular datasets, while we use decision trees for scalability.
Figure 8 shows the average rank results, including TabPFN v2-SQ and TabPFN v2-DT alongside variants using bagging and KMeans-based sampling. We observe that all variants improve upon the vanilla TabPFN v2 on large-scale datasets, with TabPFN v2-DT and TabPFN v2-SQ achieving the most significant improvement.
@@ 8 Conclusion
Our analysis highlights TabPFN v2’s potential as a pre-trained foundation model for tabular data. By leveraging randomized feature tokens, it effectively handles dataset heterogeneity and excels in in-context learning, outperforming other models on small- to medium-scale datasets. Additionally, its ability to create nearly separable embeddings suggests its potential role as a feature encoder. To address TabPFN v2’s limitations on high-dimensional, large-scale, and many-category datasets, we propose post-hoc divide-and-conquer strategies that improve scalability, expanding applicability while preserving its core strengths.
Appendix A Additional Related Work
Learning with Tabular Data. Tabular data is prevalent across diverse fields, including healthcare, finance, and education (Kovalerchuk & Vityaev, 2005; Hyland et al., 2020; Romero & Ventura, 2010; Amatriain et al., 2010). Tree-based models, such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), have long dominated this domain. However, recent advances in deep neural networks (DNNs) have demonstrated strong potential for tabular data (Borisov et al., 2024). Popular architectures like multi-layer perceptrons (Gorishniy et al., 2021; Kadra et al., 2021) and Transformers (Huang et al., 2020) have been adapted to tabular tasks, alongside custom architectures designed specifically for tabular data (Klambauer et al., 2017; Wang et al., 2021).
Deep tabular methods can be broadly categorized into two types. The first type directly processes raw features (Holzmüller et al., 2024; Gorishniy et al., 2024; Ye et al., 2025), sometimes incorporating feature-specific encoding strategies (Gorishniy et al., 2022). The second type tokenizes features, transforming an example into a set of tokens (Song et al., 2019; Huang et al., 2020; Rubachev et al., 2022). Comprehensive benchmarks have been developed to evaluate these methods across diverse datasets (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025), highlighting the strengths and weaknesses of deep tabular models in various scenarios.
Appendix B Evaluation Details
@@
Please refer Ye et al. (2024b) for details of the 300 small to medium datasets. For high-dimensional datasets, we selected 18 datasets with more than 2000 features from the scikit-feature repository. Detailed statistics of high-dimensional datasets and large-scale datasets are reported in Table 3 and Table 4.
@@
We follow Ye et al. (2024b) and use different colors to represent various categories of methods in the result figures, ensuring clarity and easy comparison. In Figure 2 and Figure 3, we compare the following methods:
@@
  • Classical Methods ( ): The classical methods include Dummy, Logistic Regression (LR), K-Nearest Neighbors (KNN), Support Vector Machines (SVM), Naive Bayes, Linear Regression, and DNNR (Nader et al., 2022), which serve as basic baselines for classification and regression tasks.
@@
  • Tree-based Methods ( ): Tree-based methods such as Random Forest (Breiman, 2001), XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018) are known for their high performance on tabular data.
@@
  • MLP variants ( ): MLP variants, including vanilla MLP, MLP-PLR, Self-Normalizing Neural Networks (Klambauer et al., 2017), and Residual Network (Gorishniy et al., 2021), enhance the flexibility and generalization of traditional MLP architectures through advanced regularization and residual connections.
@@
  • Special Architectures ( ): Methods with specially designed architectures, such as DCNv2 (Wang et al., 2021), DANets (Chen et al., 2022), and TabCaps (Chen et al., 2023), focus on improving feature interaction and abstraction to capture complex relationships in tabular data.
@@
  • Token-based Methods ( ): Token-based methods like AutoInt (Song et al., 2019), TabTransformer (Huang et al., 2020), FT-Transformer (Gorishniy et al., 2021), and ExcelFormer (Chen et al., 2024) represent features as tokens, enabling models to capture higher-order interactions through attention mechanisms.
@@
  • Regularization-based Methods ( ): Regularization-based methods, including TANGOS (Jeffares et al., 2023), SwitchTab (Wu et al., 2024), and PTaRL (Ye et al., 2024a), aim to improve model generalization by incorporating regularization techniques during training to enhance the robustness of predictions.
@@
  • Tree-mimic Methods ( ): Tree-mimic methods, such as NODE (Popov et al., 2020), GrowNet (Badirli et al., 2020), and TabNet (Arik & Pfister, 2021), combine the interpretability of decision trees with the power of deep learning, employing attention mechanisms to select important features.
@@
  • Context-based Methods ( ): Context-based methods like TabR (Gorishniy et al., 2024) and ModernNCA (Ye et al., 2025) leverage contextual information from the training data to improve predictions by utilizing neighborhood-based and in-context learning strategies.

### v2

@@ 6.2 Leave-one-fold-out Feature Extraction
To address this challenge, we propose a leave-one-fold-out strategy that enables the extraction of comparable embeddings for training and test data. In the TabPFN v2 framework, we treat examples with true labels as the support set ${\mathcal{S}}$, and those with dummy labels as the query set ${\mathcal{Q}}$. Under the standard configuration, ${\mathcal{S}}$ corresponds to the labeled training set and ${\mathcal{Q}}$ to the unlabeled test instances. To extract comparable embeddings for the training examples, they must also be included in ${\mathcal{Q}}$ with dummy label embeddings. This, however, creates a dilemma: effective in-context learning relies on maximizing the size of ${\mathcal{S}}$ to ensure sufficient knowledge transfer to ${\mathcal{Q}}$. Including training examples in ${\mathcal{Q}}$ thus competes with the need to keep ${\mathcal{S}}$ as large as possible.
To overcome this dilemma, we partition the training set into multiple folds (e.g., 10). In each round, one fold serves as ${\mathcal{Q}}$—with dummy labels used for embedding extraction—while the remaining folds form ${\mathcal{S}}$ with true labels. This setup preserves sufficient label supervision in ${\mathcal{S}}$ while enabling the extraction of embeddings for training instances in ${\mathcal{Q}}$. Results in Figure 5 (c)-(f) show that embeddings extracted by this strategy (with 10 folds) more faithfully capture dataset structure. We observe that TabPFN v2 simplifies the original tabular data distributions, transforming datasets into nearly linearly separable embedding spaces—especially after intermediate transformer layers.
6.3 Validation of Embedding Quality
To validate the quality of the extracted embeddings, we train a logistic regression on embeddings derived from the training set and evaluate it on test set embeddings.
Table 1: Average rank (lower is better) of TabPFN v2 and linear classifiers trained on the extracted embeddings across 29 classification datasets. Combined: embeddings from up to three layers (from the 12 available layers) are selected and concatenated, based on the validation set performance.
@@ 6.3 Validation of Embedding Quality
Rank 2.69 5.97 4.28 4.00 2.12 1.94
The average rank across 29 classification datasets from the tiny benchmark2 in [76] is reported in Table 1. Remarkably, training a simple linear classifier on the extracted embeddings achieves performance comparable to that of TabPFN v2’s in-context learner. Furthermore, concatenating embeddings from multiple layers (e.g., both output and intermediate representations) can sometimes lead to even better results. These findings underscore TabPFN v2’s potential as a strong and versatile feature encoder, suggesting broader applicability in downstream tasks such as tabular data analysis.
7 Improving TabPFN v2 via Test-Time Divide-and-Conquer
This section addresses the limitations discussed in Section 4.3, aiming to extend TabPFN v2’s applicability beyond the boundaries. Specifically, we propose post-hoc divide-and-conquer strategies inspired by Chain-of-Thought (CoT) prompting [71], which decompose challenging tasks into simpler subtasks that TabPFN v2 can effectively handle.
@@ 7.1 High Dimension Datasets
High-dimensional datasets [37] present a unique challenge due to the quadratic complexity of TabPFN v2 with respect to the number of dimensions. To mitigate this, we propose subsampling the feature space into smaller subsets, processing each subset independently, and combining the predictions in an ensemble (bagging) fashion, similar to random forests [8].
In detail, we iteratively sample $m$ subsets, each containing $d^{\prime}<d$ randomly selected attributes. For each subset, we leverage TabPFN v2’s ability to handle lower-dimensional data to obtain predictions. We denote this divide-and-conquer and then ensemble strategy as TabPFN v2, which aggregates outputs using averaging (for regression) or majority voting (for classification).
Figure 6 (left) summarizes the results on 18 high-dimensional classification datasets. A variant that utilizes PCA to reduce the dimensionality, together with bagging, resolves the dimensionality issue to some extent. TabPFN v2 with $d^{\prime}=500$ and $m=\left\lceil d/{d^{\prime}}\right\rceil$ significantly increases the mean accuracy (to the highest), effectively extending TabPFN v2’s scalability to datasets with $d\geq 2000$.
@@ 7.2 Multi-Class Problems with More Than 10 Classes
For a task with $C>10$ classes, we encode each label $y\in[C]$ as a $t$-digit decimal representation, where $t=\lceil\log_{10}C\rceil$. For each digit position $j\in\{1,\ldots,t\}$, we train a separate TabPFN v2 model $f_{j}$ to predict the $j$-th digit. During inference, the predicted digits are reconstructed to obtain the final class label. This strategy is also developed in [49], and we denote it as TabPFN v2-DPT. As the decimal encoding inherently introduces artificial correlations among classes — classes that share the same digit at a given position are grouped together, even if they are semantically unrelated. To mitigate this effect, our TabPFN v2 randomly permutes the class-to-digit mapping $\sqrt{C}$ times, leading to different groupings in each run, and the prediction results are ensembles to improve robustness. As shown in Figure 6 (middle), this approach achieves the second-best mean accuracy on 12 datasets with more than 10 classes while preserving computational efficiency.
Figure 6: “*” indicates our extension. Left: Mean accuracy on 18 high-dimensional datasets. “-PCA” is another variant using PCA to reduce dimensions. Middle: Mean accuracy on 12 datasets with more than 10 classes. “-ECOC” denotes the multi-class ECOC strategy implemented by [29]. Right: Average rank on 18 large-scale datasets. “-B” refers to the variant that randomly subsamples 10,000 training examples four times and aggregates their predictions. “-K” denotes the variant that selects a representative subset of 10,000 training examples based on proximity to prototypes obtained via KMeans. All variants improve TabPFN v2. Detailed results can be found in Appendix E.
@@ 7.3 Large-Scale Datasets
For large-scale datasets, we randomly sample 10,000 training examples from the full training set as the support set and treat the remaining training examples and test instances as the query set. We extract their embeddings to form a new tabular dataset, on which a logistic regression classifier is trained to make predictions on the test set embeddings. This process is repeated four times, and the final predictions are aggregated. We denote this version as TabPFN v2-SQ.
We also investigate integrating TabPFN v2 with decision trees to handle large-scale tasks. We note that a similar strategy was mentioned in [29] to handle within-dataset heterogeneity for a drastically different purpose. Specifically, we sample 32 subsets from the original training set, each containing 60% of the original data (sampled without replacement). For each subset, we first train a shallow decision tree by setting the minimum number of samples required to split an internal node to 10,000. The decision tree partitions the training set into smaller, more manageable subsets. During inference, a test instance is first passed through each of the shallow decision tree to a leaf node and then predicted by the corresponding TabPFN v2 model. The predictions from all 32 models are aggregated. We denote this extension as TabPFN v2-DF.
Figure 6 (right) shows the average rank results, including TabPFN v2-SQ and TabPFN v2-DF alongside variants using bagging and KMeans-based sampling. We observe that all variants improve upon the vanilla TabPFN v2 on large-scale datasets, with TabPFN v2-DF and TabPFN v2-SQ achieving the most significant improvement.
@@ 8 Conclusion
We present a timely investigation into TabPFN v2, a groundbreaking foundation model for tabular tasks. Our analysis uncovers the core mechanism behind TabPFN v2’s strong performance across heterogeneous tabular datasets: it can infer attribute relationships on-the-fly—even from randomly initialized token inputs—without relying on pre-defined semantics or learning dataset-specific representations. We also demonstrate that TabPFN v2 can be repurposed as a powerful feature encoder, enabling broader applications such as data visualization and diagnostic analysis. To address its limitations in more complex data regimes, we introduce post-hoc divide-and-conquer strategies that extend TabPFN v2’s utility without requiring model re-training. Together, these contributions offer fresh insights into advancing the development and application of foundation models for tabular data.
Appendix A Evaluation Details
Experimental Compute Resources. All experiments were conducted using 4 NVIDIA RTX 6000 Ada GPUs and 2 Intel(R) Xeon(R) Platinum 8352V CPUs.
@@ Appendix A Evaluation Details
Please refer [76, 46] for details of the 300 small to medium datasets. For high-dimensional datasets, we selected 18 datasets with more than 2000 features from the scikit-feature repository. Detailed statistics of high-dimensional datasets and large-scale datasets are reported in Table 2 and Table 3. The abbreviations of feature names shown in Figure 4 are explained in Tables 4 and 5.
@@
We follow [76] and use different colors to represent various categories of methods in the result figures, ensuring clarity and easy comparison. In Figure 1 (right) and Figure 3, we compare the following methods:
@@
  • Classical Methods ( ): The classical methods include Dummy, Logistic Regression (LR), K-Nearest Neighbors (KNN), Support Vector Machines (SVM), Naive Bayes, Linear Regression, and DNNR [53], which serve as basic baselines for classification and regression tasks.
@@
  • Tree-based Methods ( ): Tree-based methods such as Random Forest [8], XGBoost [12], LightGBM [39], and CatBoost [57] are known for their high performance on tabular data.
@@
  • MLP variants ( ): MLP variants, including vanilla MLP, MLP-PLR, Self-Normalizing Neural Networks [42], Residual Network [23], and TabM [20] enhance the flexibility and generalization of traditional MLP architectures through advanced regularization and residual connections.
@@
  • Special Architectures ( ): Methods with specially designed architectures, such as DCNv2 [69], DANets [10], and TabCaps [9], focus on improving feature interaction and abstraction to capture complex relationships in tabular data.
@@
  • Token-based Methods ( ): Token-based methods like AutoInt [65], TabTransformer [32], FT-Transformer [23], and ExcelFormer [11] represent features as tokens, enabling models to capture higher-order interactions through attention mechanisms.
@@
  • Regularization-based Methods ( ): Regularization-based methods, including TANGOS [36], SwitchTab [73], and PTaRL [79], aim to improve model generalization by incorporating regularization techniques during training to enhance the robustness of predictions.
@@
  • Tree-mimic Methods ( ): Tree-mimic methods, such as NODE [56], GrowNet [5], and TabNet [4], combine the interpretability of decision trees with the power of deep learning, employing attention mechanisms to select important features.
@@
  • Context-based Methods ( ): Context-based methods like TabR [22] and ModernNCA [77] leverage contextual information from the training data to improve predictions by utilizing neighborhood-based and in-context learning strategies.
@@
Remark. The standard checkpoint released by TabPFN employs a feature grouping size of 2, which complicates the analysis of individual feature embeddings and inter-feature relationships. To facilitate such analysis, we use a modified checkpoint with group size=1 for the experiments in Figure 4, which is available at HuggingFace.